OpenStudy (anonymous):

Solve (∂^2u/∂x^2) + (∂^2u/∂y^2) = 0 in y > 0 and −∞ < x < ∞ with u(x, 0) = f(x).

OpenStudy (anonymous):

Is this Calc or Diff eq?

OpenStudy (anonymous):

You still there?

OpenStudy (anonymous):

I'm willing to work through this with you. A reply would be nice.

OpenStudy (anonymous):

It's differential equations, it's for a mathematical physics class

OpenStudy (anonymous):

the classic example of \(\displaystyle \frac{\partial^2 u }{\partial x^2}+\frac{\partial^2 u }{\partial y^2}=0\) is \(\displaystyle u(x,y)=e^x \cos y + e^y \sin y\)

OpenStudy (anonymous):

Sorry, I should specify that we are using fourier transforms to solve these problems